# Towards AI-Architecture Liberty: A Comprehensive Survey on Designing and Collaborating Virtual Architecture by Deep Learning in the Metaverse

### v3

@@ 2.3.5. Generative Framework
Furthermore, some studies have shown promise within generative frameworks in connecting 3D geometries and algorithmic manipulation, as well as merging geometry utilizing latent space (Miguel et al., 2019; Sebestyen et al., 2021). This exploration involves investigating hyperparameters and loss functions in GANs while variable manipulation in the latent space of Variational Autoencoders (VAEs). For instance, Mueller examined various hyperparameter options and combinations in GANs to understand their impact on generating meaningful single-family home buildings (Mueller, 2023). In contrast, Sebestyen et al. delved VAE deeper into manipulating geometry by adjusting variables in latent space and incorporating semantic information to achieve intentional output results (Sebestyen et al., 2021).
@@ 3. Generated 3D Architecture: A PARADIGM SHIFT
In this section, we investigate the generative approaches of virtual architecture utilizing various DGMs across the fields of CV (Fig.  8, Table  3 and Table  3.4) and architecture (Table  4). Firstly, we introduce some research on architectural designs concentrating on 2D deep generative models, particularly GANs, with the goal of 3D transposition (Section 3.1). Subsequently, we categorize deep learning generation approaches for 3D representations into four groups based on DGMs (Section 3.1), including GAN (Section 3.2), VAE (Section 3.3), 3D-aware image synthesis (Section 3.4), and diffusion model (Section 3.5).
Figure 8. A systematic taxonomy for a review of generation approaches on virtual architecture design with DGMs.
3.1. 3D Form Transposition with Constrained Approaches
In the past, DGMs have primarily focused on 2D image processing tasks, such as texture synthesis (Gao et al., 2021), style transfer (Karras et al., 2019), photorealistic image generation (Nichol et al., 2022a), and text-to-image synthesis (Nichol et al., 2022a). Architectural research has similarly centered around 2D to 3D transposition using image-to-image translation networks like Style Transfer (Özel, 2020; Ren and Zheng, 2020; Liu et al., 2021; Zhang, 2019), StyleGAN (del Campo et al., 2019; Zhang, 2019; Zhang and Huang, 2021; Karras et al., 2019), and pix2pixGAN (Yu, 2020; Di Carlo et al., 2022).
@@
This approach converts 3D representations into generative 2D images through computational calculations or manual adjustments. The process starts by segmenting a 3D model into discrete images, such as sections, plans, and projections from various viewpoints. For example, Zhang and Blasetti combined two architectural forms using generated 2D-pixel images through discrete section slicing and multiple views (Zhang, 2019).
Some research employs 2D manipulation of spatial features on the plan or section, drawing from traditional architectural design principles. This method offers easy usability and high efficiency in terms of time and computation. For instance, EI Asmar and Sareen utilized latent space through interpolations to generate various images and further generate 3D voxelization by vector arithmetic (Asmar and Sareen, 2020). Similarly, Del Campo et al. also utilized the latent space of StyleGAN to perform feature interpolation to form the continuous sequence of images through latent walks (del Campo et al., 2022). Based on the above method, Bank et al. further developed an interactive tool that can manipulate data in latent spaces, where the generated stylized images are represented as point clouds and can be assembled as spectral entities in adjustable resolution (Bank et al., 2022). Their research significantly advances by using pixel projection to convert values in pixels into a 3D model. In contrast, other approaches utilize latent space to store downscaled perspective information (using a 2D image to store 3D information) for further manipulation. Huang et al. employed latent space to encode the 3D information from images of GANs (Huang et al., 2021). They conduct graph sampling of the generative images and interpolate the rendering image results with a sequence of perspectives leveraging latent space. The perspectival GAN is an extended research comprising latent space rotation to learn 3D information in 2D images (Kim and Huang, 2022). This reflects the scale interpretations from various perspectives in DL-assisted generation as a critical focus on spatial perception following the architecture principle.
The 3D form transposition is evidence of computational freedom in the innovative 3D generation since the future direction of virtual spaces lies in complex variations. This generation method allows for the preservation of high resolution in both input and output, as 2D images are lightweight and simple to process. Furthermore, training algorithms for 2D-based GAN networks have been well-developed, thereby providing a wide range of possibilities for human interaction with algorithms, as well as various adjustable output options through contouring 2D patterns into 3D forms.
3.2. 3D Solid Form Generation with the GANs
@@
GANs generate 3D form through training distributions by playing zero-sum games with uncertain but high-quality output. Thus, it leads to typical applications in testing the agencies of AI, opting for conceptual design, and democratizing design.
Method Names
Publication& Year
3D Representation Deep Generative Model
3D GAN (Wu et al., 2016)
NIPS 2016
Voxel grid
GAN
Text2Shape (Chen et al., 2019)
ACCV 2018
Voxel grid
GAN
PLATONICGAN (Henzler et al., 2019)
CVPR 2019
Voxel grid
GAN
IG GAN  (Lunz et al., 2020)
arXiv 2020
Voxel grid
GAN
Achlioptas et al.  (Achlioptas et al., 2018)
ICML 2018
Point cloud
GAN
Shu et al.  (Shu et al., 2019)
CVPR 2019
Point cloud
GAN
Get3d  (Gao et al., 2022)
NeurIPS 2022
Mesh
GAN
IM-Net  (Chen and Zhang, 2019)
CVPR 2019
Neural field
GAN
Kleineberg et al.  (Kleineberg et al., 2020)
arXiv 2020
Neural field
GAN
Brock et al.  (Brock et al., 2016)
arXiv 2016
Voxel grid
VAE
Autosdf  (Mittal et al., 2022)
CVPR 2022
Voxel gird
VAE
Sagnet  (Wu et al., 2019)
TOG 2019
Voxel gird
VAE
Li et al.  (Li et al., 2020)
AAAI 2020
Voxel gird
VAE
Pq-net  (Wu et al., 2020)
CVPR 2020
Voxel gird
VAE
AdversarialAE  (Zamorski et al., 2020)
CVPR 2020
Voxel gird
VAE
Multi-Chart  (Ben-Hamu et al., 2018)
TOG 2018
Mesh
VAE
SDM-NET  (Gao et al., 2019)
TOG 2019
Mesh
VAE
Tm-net  (Gao et al., 2021)
TOG 2021
Mesh
VAE
Polygen  (Nash et al., 2020)
ICML 2020
Mesh
VAE
PointFLow  (Yang et al., 2019)
CVPR 2019
Point cloud
Normaliz. flow model
CLIP-Forge  (Sanghi et al., 2022)
CVPR 2022
Voxel grid
Normaliz. flow model
DreamFusion (Poole et al., 2022)
arXiv 2022
Mesh
Diffusion model
PVD  (Zhou et al., 2021)
CVPR 2021
Hybrid: point cloud-voxel
Diffusion model
Magic3D  (Lin et al., 2023)
arXiv 2022
Neural field-Mesh
Diffusion model
LION  (Zeng et al., 2022)
arXiv 2022
Mesh
Diffusion model
Point-E  (Nichol et al., 2022b)
arXiv 2022
Neural field-Point cloud
Diffusion model
Table 3. An overview of 3D generative approaches of 3D shape generation. For 3D shape generation, each method allows the generating of editable models for explicit representations.
@@
3.2.2. 3D Solid Form Generation
(del Campo et al., 2019)
2D to 3D
Test AI agency in design
Utilizing Style Transfer to train two datasets Baroque and Modern images as a basis to form a 3D model
(Özel, 2020)
2D to 3D
HCI in urban design
Utilizing Style Transfer to generate different stylized images and generate 3D geometry through procedural modeling
(Ren and Zheng, 2020)
2D to 3D
Test AI agency in design
Utilizing Style Transfer to replace pixels with voxelization units to generate 3D forms
(Liu et al., 2021)
2D to 3D
Toolkits for 3D generation
Utilizing Style Transfer to assist the generation of 3D structure from 2D images
(Yu, 2020)
2D to 3D
Generate building massing
Utilizing pix2pixGAN to generate plan pattern and section pattern, then converted to 3D massing
(Di Carlo et al., 2022)
2D to 3D
Generate building massing
Utilizing pix2pixGAN to generate urban morphology to create building massing
(Zhang, 2019)
2D to 3D
Form finding to assist design
Slicing a 3D model and trained with different combinations of 2D styleGAN networks, and finally stitching into a 3D model
(Zhang, 2019)
2D to 3D
Form finding to assist design
3D model generation based on 2D plan and section using Style Transfer
(Zhang and Huang, 2021)
2D to 3D
Form finding to assist design
Combining the spatial sequence information to generate 3D form from 2D images through multi-level deep generative networks such as styleGAN
(Bank et al., 2022)
2D to 3D
Human and neural network interface
Utilizing 3D solid for training to map spatial semantics to a latent space assembled using point cloud representations
Point cloud
GAN
(Asmar and Sareen, 2020)
2D to 3D
Integrate latent space in design
GAN allows for navigation in the latent space to create digital designs using vector arithmetic and interpolation techniques, then converting resulting images to 3D voxel structures
Voxel grid
GAN
(Huang et al., 2021)
2D to 3D
Recognize 2D pattern to 3D form
Utilizing Latent space rotation and perspective projection to generate 3D model
Voxel grid
GAN
(Kim and Huang, 2022)
2D to 3D
Recognize 2D pattern to 3D form
Utilizing Latent space rotation and perspective projection to generate 3D model
Voxel grid
GAN
(Çakmak, 2022)
3D Solid
Extend design cognition
Utilizing a GAN Model with a pair of encoder-decoder to process datasets and generate new 3D models, resulting in different 3D representations like point cloud and mesh
Point cloud/Mesh
GAN (with Autoencoder)
(Ondrej Vesel‘y, 2022)
3D Solid
Generate building massing
Utilizing 3D BAG dataset as a basis to train urban morphology data to generate 3D building massing
GAN
(Mueller, 2023)
3D Solid
Generate building massing
Utilizing BuildingNet dataset (Selvaraju et al., 2021) as a basis to train urban morphology data to generate 3D building massing
GAN
(Miguel et al., 2019)
3D Solid
Generation, manipulation and form finding of structural typologies
Utilizing VAE to learn continuous latent space to generate new geometries
Voxel grid
VAE
(Kahraman et al., 2021)
3D Solid
Solve design problems incorporating deep learning
Utilizing neural network workflows (MoVAE and MoGAN) to solve complex, multi-objective design problems
Voxel grid
Base VAE & GAN
(Sebestyen et al., 2021)
3D Solid
New way to design models
Utilizing VAE to encode and decode geometries through dimensionality manipulation
Voxel grid
VAE
(Zhang, 2020)
NLP-3D Solid
Semantic dataset assisted design
Utilizing language model to predict housing plan by training large dataset relating texts to forms
(Guida, 2023)
NLP-3D Solid
Multimodal generation with text input
Utilizing diffusion models to generate 3D forms from text input with multimodality
Diffusion model
Table 4. The related works in the architecture fields are categorized into three types: 2D to 3D transposition, 3D solid generation, and NLP-based 3D form generation. The ‘Category’ column represents the generation approaches: ‘2D to 3D’ means generating 3D form based on 2D images; ‘3D Solid’ means generating 3D form directly with DGMs including GAN, VAE, and diffusion model. ‘NLP’ means the 3D form generation process includes text input to assist the output control. The ‘Objective’ column represents the directions and objectives the research aims to address. The ‘Methodology’ column provides an overview of the workflow proposed in the generation process. ‘3D Representation’ indicates generated data formats, encompassing point cloud, voxel grid, and mesh. ‘Generative Model’ indicates the deep generative model type employed in a work.
This refers to generating 3D solid form through a direct 3D data acquisition, evaluation, transformation, and rearrangement using deep generative models  (Steinfeld et al., 2019). Currently, architectural design that utilizes GANs in 3D solid form generation is based on explicit representations, including voxel grids, point clouds, and meshes. This approach, aligning training datasets with spatial design parameters, is widely applied in architectural design where GANs, as the generative framework, frequently achieve specific design objectives.
As mentioned, GANs can generate 3D buildings with representations such as mesh and point cloud from a dataset with labels or semantic interpretation. For instance, Immanuel Koh uses a 3D GAN network to generate innovative housing typologies automatically through training a large dataset of both exterior and interior Singapore high-rise buildings (Koh, 2022). This work inherits the configurations of architectural forms by extracting building block arrangements. Another work follows an architectural design strategy based on modeling rules. Zheng and Yuan used vector-based neural networks to find forms based on NURBS  [^13] Non-uniform rational basis spline, a mathematical model using basis splines to represent complex curves, surfaces, and solid form. rules with different design features (Zheng and Yuan, 2021).
Moreover, integrating additional computational methods or generative models with 3D GANs can expand to wider design cases for 3D form generation. For instance, Puteneers uses Houdini combined with 3D GAN as a form-finding tool to test the agency of algorithms regarding creating artifacts, coupled with image recognition of 3D artifacts in a project named Ugly & Stupid  [^14] Source: https://putteneersjoris.xyz/projects/Ugly%20Stupid%20Honest/ugly_stupid_honest.html. In comparison, Cakmak added a pair of encoder-decoders in the GAN framework to process and save data for generating 3D models (Çakmak, 2022). This also aims to extend design cognition by adding AI as an agent to understand the mind in cognition science using embodied action. Beyond the studies in digital theory and design experiments, 3D GANs have also been explored as tools to conduct urban design or to address urban development challenges. A noteworthy study has been conducted on geometry extraction within urban environments using 3D GANs  (Ondrej Vesel‘y, 2022). This enables generated 3D model with GAN automatically through 3D BAG   [^15] An overview of 3D BAG, source:  https://docs.3dbag.nl/en/ including different levels of details. The model was trained based on three layers of information: building geometry, site context, and area of interest. The information is stored in raster data for 3D representation of voxel grids in deep learning. For addressing urban design challenges, Mueller used 3D GAN to efficiently produce feasible building geometries, aiming to fulfill increasing demands on housing shortage (Mueller, 2023).
3.2.3. Limitations
Although GAN is the most popular and broad DGM regarding deep-learning-assisted architectural generation, these methods have several limitations, including style variation, category singleness, design unpredictability, and topological inconsistency.
Fewer variations in style. The performance of GAN frameworks heavily depends on the quality and nature of the input data, resulting in limited variations in style. In the context of 2D to 3D form finding, many studies have relied on StyleGAN as the basis for form generation, which tends to produce outputs that mimic the style of the “style image” without adjustable options.
Singleness of category. Since one training process can only process one dataset category, this output is constrained to the category for design purposes. For example, the 3D-GAN-Housing project has a certain degree of repetition in high-rise building design due to the limitation of the trained structure (Koh, 2022). Since the dataset is limited, the blocks always follow the same evolutionary rule, which lowers the variation in style and is limited to specific categories. This can be adapted to modular housing design; however, not suitable for virtual space generation.
Unpredictability in design. The image-to-image translation of GANs is characterized by unpredictability. The training process of these algorithms is time-consuming and requires substantial computing resources. Also, the generative logic underlying these processes allows designers to evaluate their effects only once they observe the final output. The latent vector undergoes arbitrary modifications in different epochs, adding to the complexity and unpredictability of the output.
Topological inconsistency. Firstly, when the 3D forms are constructed solely from 2D images, these forms will inevitably carry traces of the slicing process, leading to a loss of interior details and the overall consistency of the structure. Secondly, using a constrained 3D segmentation algorithm poses a significant challenge in generating consistent forms, leading to topological inconsistencies in the form of gaps and defects in the final output. For instance, applying this method to the reconstruction of furniture reveals inconsistencies in the generated 3D shapes, as the algorithm needed to be pre-trained on the specific type of objects (Wu et al., 2016).
Computing requirements of 3D data. Compared to other models, GAN models can typically produce 3D structures that are more detailed and realistic. However, they are also more unstable and challenging to train. Converting data from 2D to 3D usually takes a long time.
3.3. Architectural Form from Latent Space with Variational Autoencoder
VAE results in an accurate 3D generation but lower resolution, compared to GANs’ uncertain while high-quality output. Thus, its applications are commonly generated for different criteria and scenarios against visual output. For instance, Azizi et al. generated reliable and plausible architectural compositions by VAE through encoding and decoding information about people’s spatial movements and activities (Azizi et al., 2020)
@@ 3.3. Architectural Form from Latent Space with Variational Autoencoder
VAEs utilize pointwise loss to find a probability density by explicit representations to obtain an optimal solution by minimizing a lower bound on the log-likelihood function. Brock et al. introduced the first variational autoencoder (VAE) for processing 3D voxel grids to address the instability issues inherent in GAN approaches (Brock et al., 2016) . This method employs a pair of encoder-decoder architecture: the encoder, comprising four 3D convolutional layers, maps information to latent vectors, while the decoder reconstructs these vectors into 3D voxels. Subsequent work has focused on enhancing the quality of voxel reconstructions, particularly in producing smooth rounded edges (Mittal et al., 2022). For point cloud representation, the challenges associated with GAN instability have led to the development of alternative generative models based on encoders. Variational autoencoders (VAEs) and adversarial autoencoder models (AAEs) have emerged as viable alternatives (Zamorski et al., 2020). Similarly, generating meshes with VAEs presents challenges akin to those faced with GANs. Given the complexity of processing topology, approaches like the multi-chart parameterization of meshes have been proposed to handle irregular mesh structures (Ben-Hamu et al., 2018). Several methods have sought to simplify this process, including SDM-Net-based approaches like TM-Net, which defines a textured space on a template cube mesh (Gao et al., 2019, 2021; Nash et al., 2020).
3.3.2. Architectural Information Extraction from Latent Space

### v4

@@ 2.3.5. Generative Framework
Furthermore, some studies have shown promise within generative frameworks in connecting 3D geometries and algorithmic manipulation, as well as merging geometry utilizing latent space (Miguel et al., 2019; Sebestyen et al., 2021). This exploration involves investigating hyperparameters and loss functions in GANs while variable manipulation in the latent space of Variational Autoencoders (VAEs). For instance, Mueller examined various hyperparameter options and combinations in GANs to understand their impact on generating meaningful single-family home buildings (Mueller, 2023). In contrast, Sebestyen et al. delved a VAE deeper into manipulating geometry by adjusting variables in latent space and incorporating semantic information to achieve intentional output results (Sebestyen et al., 2021).
@@ 3. Generated 3D Architecture: A Paradigm Shift
To examine the gap between the two fields and the potential advantages of each approach for virtual architecture, we investigate various generative approaches using different DGMs across the fields of CV (Fig.  8) (Table 3 and Table A, Appendix) and architecture (Table  5, Appendix). Since the existing related architectural works do not systematically form solid design philosophies, we elaborate on these works following their emphasized generative approaches, including 3D transposition (Section 3.1), generation by GANs (Section 3.2), VAEs (Section 3.3), 3D-aware image synthesis (Section 3.4), and diffusion models (Section 3.5). The first approach relies on additional transposition from generative images, while the others generate geometry directly.
Figure 8. A systematic taxonomy for reviewing generation approaches on virtual architecture design with DGMs.
3.1. Constrained Approaches Using 3D Form Transposition
In the past, DGMs have primarily focused on 2D image processing tasks, such as texture synthesis (Gao et al., 2021), style transfer (Karras et al., 2019), photorealistic image generation (Nichol et al., 2022a), and text-to-image synthesis (Nichol et al., 2022a). Architectural research has similarly centered around 2D to 3D transposition using image-to-image translation networks like Style Transfer (Özel, 2020; Ren and Zheng, 2020; Liu et al., 2021; Zhang, 2019), StyleGAN (del Campo et al., 2019; Zhang, 2019; Zhang and Huang, 2021; Karras et al., 2019), Pix2Pix GAN (Yu, 2020; Di Carlo et al., 2022), and CycleGAN (Dong et al., 2023).
@@ 3.1. Constrained Approaches Using 3D Form Transposition
This approach converts generative images into 3D representations through computational calculations or manual adjustments. The process starts by segmenting a 3D model into discrete images, such as sections, plans, and projections from various viewpoints. For example, Zhang and Blasetti combined two architectural forms using generated 2D-pixel images through discrete section slicing and multiple views (Zhang, 2019). Compared to direct 3D approaches, this method offers easier usability and higher computational efficiency as it is based on generative images.
Research in this approach utilize traditional architectural design principles to manipulate spatial features in 2D images on the plan (Ondrej Vesel‘y, 2022; Asmar and Sareen, 2020), section (Bank et al., 2022), facade (del Campo, 2022; del Campo et al., 2022), and perspective images (Huang et al., 2021; Kim and Huang, 2022). For instance, Asmar and Sareen employed a StyleGAN to leverage latent space for image generation, followed by 3D voxelization using vector arithmetic, with a specific focus on generating a single building (Asmar and Sareen, 2020). Vesely also utilized Pix2Pix GANs while generating building massings within urban environments by training on high-fidelity 3D building models  [^14] The dataset is 3D BAD. An overview of 3D BAG, source:  https://docs.3dbag.nl/en/, and incorporating multi-layered site context information (Ondrej Vesel‘y, 2022). Whereas Bank et al. experimentally developed building forms with 3D point clouds from sampling generative images aiming for teaching methods (Bank et al., 2022). In contrast, some research utilize pixel projection to convert 2D-pixel values, generated by latent walks, into a corresponding 3D model. (del Campo et al., 2022; del Campo, 2022). In these works, Del Campo et al. innovatively generated a house emancipating from a canonical AI-generative approach by intentionally challenging curve-fitting techniques and embracing data scarcity or overfitting (del Campo, 2022). In contrast to prior approaches, some researchers utilize latent space for downscaled perspective information storage, condensing 3D data into 2D images for manipulation. For example, Huang et al. generated rendering results with various perspectives by interpolating images generated from the latent space of GAN (Huang et al., 2021). This approach extends prior research by incorporating a rotation of latent space for 3D perspective drawing (Kim and Huang, 2022). This reflects a critical focus on spatial perception in DL-assisted generation, following architectural principles, and interpreting scale from various perspectives.
The 3D transposition approach exemplifies how freedom of computation and extensive manipulation shape the future of virtual architecture with its focus on intricate variations. Given that generative images are lightweight and easy to process, this method ensures high-resolution retention in both input and output. However, this method sacrifices geometry controllability and production efficiency for a generation. Additionally, significant human labor is required to establish an automatic workflow. These shortcomings have urged a paradigm shift in generative approaches towards direct geometry production.
3.2. Approaches Using GANs
@@
GANs generate 3D form through training distributions by playing zero-sum games with uncertain but high-quality output. Thus, they lead to typical applications in testing the agencies of AI, opting for conceptual design, and democratizing design.
@@
3.2.2. 3D GAN Utilized in Architecture
This refers to generating 3D solid form through a direct 3D data acquisition, evaluation, transformation, and rearrangement using deep generative models  (Steinfeld et al., 2019). This approach, aligning training datasets with spatial design parameters, usually achieves specific design objectives. Specifically, designers utilize datasets with design meaning by labels or semantic interpretation. For instance, Immanuel Koh employed 3D GAN to generate massive Singapore high-rise buildings by labeling building exteriors and interiors of training models (Koh, 2022). Additionally, researchers and designers have fully probed dataset forms related to architectural structures to address the deficiency in generating precise geometries. Zheng and Yuan innovated training coordinate vectors of building surfaces in neural networks to find forms based on NURBS  [^15] Non-uniform rational basis spline, a mathematical model using basis splines to represent complex curves, surfaces, and solid form. with different design features (Zheng and Yuan, 2021).
Innovating generative methods using 3D GANs for fidelity and accuracy remains a novel and timely topic in AI-assisted architectural design. Some researchers integrate additional computational methods or generative models with 3D GANs to expand the range of design cases for 3D form generation. For instance, Puteneers combined Houdini with 3D GANs as a form-finding tool to test the algorithms’ capabilities in creating artifacts as AI agency  [^16] Source: https://putteneersjoris.xyz/projects/Ugly%20Stupid%20Honest/ugly_stupid_honest.html. Similarly, Cakmak added a pair of encoder-decoders to a GAN framework to process and save data for generating 3D models (Çakmak, 2022). This approach aims to extend design cognition by incorporating AI as an agent to understand the mind in cognitive science through embodied action.
Researchers have also increased their focus on training and deployment techniques for design purposes. For instance, Mueller developed a new GAN architecture capable of generating building geometry with incorporated structural features in the dataset (Mueller, 2023). This work emphasizes testing different inputs for the generator and discriminator, analyzing the training process, and exploring various hyperparameters. Such progress demonstrates a commitment to autonomously generating building geometry using 3D GANs, leveraging automation and advanced technologies to address architectural and urban design challenges.
Although GANs are widely recognized as the primary tools for deep-learning-assisted architectural generation, they exhibit limitations in architectural design applications, including issues with style variation, category singleness, design unpredictability, and topological inconsistency. These challenges drive designers and researchers to explore diverse alternative approaches in architectural design.
3.3. Approaches Using Latent Space of VAEs
VAEs result in an accurate 3D generation but lower resolution, compared to GANs’ uncertain while high-quality output. Thus, their applications are commonly generated for different criteria and scenarios against visual output. For instance, Azizi et al. generated reliable and plausible architectural compositions by a VAE through encoding and decoding information about people’s spatial movements and activities (Azizi et al., 2020).
@@ 3.3. Approaches Using Latent Space of VAEs
VAEs utilize pointwise loss to find a probability density by explicit representations to obtain an optimal solution by minimizing a lower bound on the log-likelihood function. Brock et al. pioneered the first VAE for processing 3D voxel grids to address the instability issues inherent in GAN approaches (Brock et al., 2016). This method employs a pair of encoder-decoder architecture: the encoder, comprising four 3D convolutional layers, maps information to latent vectors, while the decoder reconstructs these vectors into 3D voxels. Subsequent work has focused on enhancing the quality of voxel reconstructions, particularly in producing smooth rounded edges (Mittal et al., 2022). For point cloud representation, the challenges associated with GANs’ instability have prompted the development of alternative generative models based on encoders. Variational autoencoders (VAEs) and adversarial autoencoder models (AAEs) have emerged as viable alternatives (Zamorski et al., 2020). Similarly, generating meshes with VAEs presents challenges akin to those faced with GANs. Given the complexity of processing topology, approaches like the multi-chart parameterization of meshes have been proposed to handle irregular mesh structures (Ben-Hamu et al., 2018). Several methods have sought to simplify this process, including SDM-Net-based approaches like TM-Net, which defines a textured space on a template cube mesh (Gao et al., 2019, 2021; Nash et al., 2020).
3.3.2. VAEs Utilized in Architecture